OpenStudy (anonymous):

I have the answer can someone please check me the volume of a cube is decreasing at a rate of 0.24 ft^3/min. What is the rate of change of the side length when the side length are 2 feet.

jimthompson5910 (jim_thompson5910):

V = s^3 dV/dt = 3s^2*ds/dt -0.24 = 3*2^2*ds/dt -0.24 = 12*ds/dt keep going to solve for ds/dt

OpenStudy (anonymous):

I got -12.24

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

or -.02

jimthompson5910 (jim_thompson5910):

-0.24 = 12*ds/dt -0.24/12 = ds/dt -0.02 = ds/dt ds/dt = -0.02

OpenStudy (anonymous):

ok I subtracted instead of divided

jimthompson5910 (jim_thompson5910):

so the side is changing at a rate of -0.02 ft/min when s = 2 ft

OpenStudy (anonymous):

so the .02 is the rate of change

jimthompson5910 (jim_thompson5910):

yes a decrease of 0.02 ft/min

OpenStudy (anonymous):

what do you mean decrease

jimthompson5910 (jim_thompson5910):

I guess I should say that the side is decreasing at a rate of 0.02 ft/min when the side length is 2 ft

OpenStudy (anonymous):

do you mean -.02

jimthompson5910 (jim_thompson5910):

rate of change = -0.02 ft/min is the same as saying "the side is decreasing at a rate of 0.02 ft/min" so yes

OpenStudy (mertsj):

decrease means gets smaller